Grade 11Organic Chemistry

the density of air at 0 degree celcius and 760mm hg pressure is 1.29g/lfind the density at at 30 degree celcius and 750 mm hg pressure

hussain cm
14 Years agoGrade 11

1 Answer

Chetan Mandayam Nayakar
14 Years ago

PV=nRT,n/V=P/RT,nM/V=PM/RT=density